A bag contains a collection of distinguishable marbles. The bag has three red marbles, six green ones, one lavender one, five yellows, and two orange marbles. HINT [See Example 7.] How many sets of four marbles include exactly two green marbles? (Note that this means there must be two other non-green marbles as well).
Answer #1

sets of four marbles include exactly two green marbles =N(select 2 from 6 marbles and 2 from 11 marbles)=(6C2)*(11C2) =15*55 =825